OpenStudy (anonymous):

Find an equation of the tangent line to the graph of f(x)=3x^3 +2x when x=1

OpenStudy (anonymous):

actually you need to take the derivative \(f'(x)\) then find \(f'(1)\) for the slope

OpenStudy (anonymous):

First find the derivative of the function: \[f' (x) = 9x^2 + 2\] Now plug in 1 instead of x: \[f'(x) = 9 + 2 = 11\] So the slope is 11, now we need to find y so we plug in 1 instead of x in the original function: \[y = 3 + 2 = 5\] Now just plug in everything to y = mx +b and find b: \[5 = 11 * 1 + b\] \[ b = -6\] So the equation is: \[y = 11x - 6\]